# Accelerating Vision Diffusion Transformers with Skip Branches

### v2

@@ 3.3 Skip-DiT: Improving Feature Smoothness
A vanilla DiT can be converted into Skip-DiT by connecting shallow blocks to deep blocks with skip branches, as shown in Figure 3 (b). Let $\bm{x}$ denote the input noise embedding, and $\bm{x}^{\prime}_{l}$ represents the output at the $l$-th layer of Skip-DiT. The architecture consists of $L$ sequential DiT blocks with skip connections. Each DiT block $\mathcal{F}_{\text{DiT}}^{l}$ at block $l$ processes the features as $\bm{x}^{\prime}=\mathcal{F}_{\text{DiT}}^{l}(\bm{x})$. The $i$-th skip branch ($i\in\{1\ldots L//2\}$) connects $i$-th block to ($L+1-i$)-th block, which can be denoted as $\mathcal{F}_{\text{Skip}}^{i}(\cdot,\cdot)$. Given output $\bm{x}^{\prime}_{i}$ from the start of the skip branch and $\bm{x}^{\prime}_{l}$ from the previous layer, the skip branch aggregates them to the input to $l$-th block as:
@@
where $\oplus$ denotes concatenation, Norm represents the layer normalization, and Linear is a linear fully-connected layer. The final output $x^{\prime}_{l}$ of Skip-DiT represents the processed noise output. Each skip branch creates a shortcut path that helps preserve and process information from earlier layers, enabling better gradient flow and feature reuse throughout the network. The combination of DiT blocks and skip branches allows the model to effectively learn the underlying noise distribution while maintaining stable training dynamics.
Similarly, we initialize a class-to-video DiT with skip branches and train it from scratch on Taichi dataset, then visualize its feature smoothness as shown in Figure 2 (w/ Skip branches). The results show Skip-DiT has a more flat feature-changing landscape at both the beginning and finalizing timesteps. This justifies our initiative to enhance DiT feature smoothness with skip connections.
3.4 Skip-Cache: Caching with Skip Branches
As we have shown better feature smoothness in Skip-DiT, we can use the feature stability and skip branch property of Skip-DiT  to implement efficient DiT caching, namely Skip-Cache. The inference process for global timestep $t$ of full inference can be expressed as follows:
$\bm{x^{\prime}}^{t}=\mathcal{F}_{\text{DiT}}^{L}(\mathcal{F}_{\text{DiT}}^{L-1}(...\mathcal{F}_{\text{DiT}}^{1}(\bm{x})))$ (7)
Consider timesteps $t-1$ and $t$ where the model generates $\bm{x}^{t-1}$ conditioned on $\bm{x}^{t}$. During this step, we cache the intermediate output from the last second layer as
$\mathcal{C}^{t}_{L-1}=x_{L-1}^{t}$ (8)
For local timestep $t-1$, with cached feature $\mathcal{C}^{t}_{l-1}$ and first skip branch $\mathcal{F}_{\text{Skip}}^{1}(\cdot,\cdot)$, the inference process can be formulated as:
$x^{\prime}_{t-1}=\mathcal{F}_{\text{DiT}}^{L}(\mathcal{F}_{\text{Skip}}^{1}(\bm{x}^{\prime t-1}_{1},\mathcal{C}^{t}_{L-1}))$ (9)
At each local timestep, only $1$-th and $L$-th blocks are executed while reusing cached features from the previous global timestep through the skip branch, significantly reducing computational overhead while maintaining generation quality. This can be extended to 1:$N$ inference pattern where $\mathcal{C}^{t}_{L-1}$ will be reused for the next $N$-1 timesteps.
@@ 4 Experiments
4.1 Implementation Details
Table 1: Class-to-video generation performance. The definition of the cache step $n$ follows that in Table 3. For the UCF101 task, videos are generated using DDIM with 50 steps, while for the remaining tasks, DDPM with 250 steps is employed. Latency and speedup are calculated on one A100 GPU, with solver 250 steps DDPM. $n=i$ indicates caching the high-level feature in $x_{t}$ for reuse during the inference of $x_{t-1},x_{t-2},\dots,x_{t-n+1}$. We highlight Baseline DiT models (without caching) in grey, and the best metrics in bold.
@@ 4.1 Implementation Details
Skip-Cache
Skip-DiT 141.30 23.78 20.62 4.32 49.21 11.92 163.03 13.55 290.05 10.02 1.00$\times$
@@
To demonstrate the remarkable effectiveness of Skip-DiT  and Skip-Cache  in video generation, we employ the classic and open-source DiT model, Latte [19], as our base model. Latte consists of spatial and temporal transformer blocks, making it suitable for both class-to-video and text-to-video tasks. Hunyuan-DiT [16], the first text-to-image DiT model with skip branches, modifies the model structure by splitting transformer blocks into encoder and decoder blocks, which are connected via long skip connections, similar to UNets. In this work, we leverage Hunyuan-DiT to investigate the effectiveness of skip connections in text-to-image tasks. Additionally, we modify the structure of Latte, following the guidance in Figure 3, to evaluate the performance of skip connections in video generation tasks and explore techniques for integrating skip connections into a pre-trained DiT model.
@@
In the class-to-video task, we conduct comprehensive experiments on four public datasets: FaceForensics [28], SkyTimelapse [39], UCF101 [34], and Taichi-HD [31]. Following the experimental settings in Latte, we extract 16-frame video clips from these datasets and resize all frames to a resolution of 256$\times$256.
For the text-to-video task, the original Latte is trained on Webvid-10M [1] and Vimeo-2M [36], comprising approximately 330k text-video pairs in total. Considering that the resolution of Webvid-10M is lower than 512$\times$512 and is predominantly used in the early stages of pre-training [20], we utilize only Vimeo-2M for training Skip-DiT. To align with Latte, we sample 330k text-video pairs from Vimeo-2M. All training data are resized to a resolution of 512$\times$512, with 16 frames per video and a frame rate of 8.
@@
For the training of class-to-video tasks, we train all instances of Skip-DiT  from scratch without any initialization. During training, we update all parameters in Skip-DiT, including skip branches. For text-to-video tasks, we propose a two-stage continual training strategy as follows:
  • Skip-branch training: The models are initialized with the weights of the original text-to-video Latte model, while the weight of skip branches are initialized randomly. During this stage, we train only the skip branches until the model can roughly generate items. This stage takes approximately one day.
  • Overall training: After fully training the skip branches, we unfreeze all other parameters and perform overall training. At this stage, Skip-DiT rapidly recovers its generation capability within approximately two days and can generate content comparable to the original Latte with an additional three days of training.
This strategy significantly reduces training costs compared to training from scratch. All our training experiments are conducted on 8 H100 GPUs, employing the video-image joint training strategy proposed in Ma et al. [19]. We find that this approach significantly enhances training stability.
@@
Following previous works [47, 18] and Latte, we evaluate text-to-video models using VBench [10], Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [43], and Structural Similarity Index Measure (SSIM) [37]. VBench is a comprehensive benchmark suite comprising 16 evaluation dimensions. PSNR is a widely used metric for assessing the quality of image reconstruction, LPIPS measures feature distances extracted from pre-trained networks, and SSIM evaluates structural information differences. All the videos generated for evaluation are sampled with 50 steps DDIM [33], which is the default setting used in Latte.
For class-to-image tasks, we evaluate the similarity between generated and real videos using Fréchet Video Distance (FVD) [35] and Fréchet Inception Distance (FID) [21], following the evaluation guidelines of StyleGAN-V [42]. Latte uses 250-step DDPM [9] as the default solver for class-to-video tasks, which we adopt for all tasks except UCF101. For UCF101, we employ 50-step DDIM [33], as it outperforms 250-step DDPM on both Latte and Skip-DiT. Table 2 highlights this phenomenon, showing our methods consistently outperform DDPM-250 under comparable throughput, except for UCF101, where DDIM performs better than 250 steps DDPM.
Implementation Details of Other Caching Methods
We compare with 4 other DiT caching methods in video and image generation: ❶ T-GATE [44] reuses self-attention in the semantics-planning phase and skips cross-attention in the fidelity-improving phase. We follow Zhao et al. [47] to split these two phases. ❷ $\Delta$-DiT identifies high similarity in deviations between feature maps and reuses them at the next timestep. While this method is originally designed for images, we extend it to video DiTs by caching only the front blocks, as we observe significant degeneration when caching the back blocks. ❸ FORA [30] reuses attention features across timesteps. ❹ PAB [47] further extends it by broadcasting cross, spatial, and temporal attention features separately. All these caching methods are performed on Latte for equal comparison with Skip-Cache  on Skip-DiT .
4.2 Main Results
Class-to-video Generation
We compare the quantitative performance of Latte and Skip-DiT on four class-to-video tasks, as shown in Table 1. Skip-DiT consistently outperforms Latte in terms of FVD scores across all tasks while achieving comparable performance in FID scores, demonstrating its strong video generation capabilities. Furthermore, we observe that Skip-Cache significantly outperforms other caching methods across most metrics, incurring only an average loss of $2.37$ in the FVD score and $0.27$ in the FID score while achieving a 1.56$\times$ speedup. In comparison, only PAB [47] achieves a speedup of more than 1.3$\times$, but at the cost of a substantial average loss of $60.78$ in FVD score and $4.48$ in FID score. Notably, in the Taichi [31] task, all other caching methods exhibit significant degradation in FVD scores($\geq$ 21.5), whereas Skip-Cache experiences only a slight loss (163.06$\rightarrow$167.54). To achieve even higher speedup on the other three class-to-video tasks, we accelerated Skip-Cache with a larger cache timesteps (N=3), resulting in a 2.19$\times$ speedup with an average loss of $6.47$ in FVD and a $0.68$ improvement in FID.
Table 2: Comparison with the faster DDIM sampler. Skip-Cache  is evaluated with a 250-steps DDPM and compared to the DDIM sampler under similar throughput. Baseline DiT models (without caching) are highlighted in grey, and the best metrics are indicated in bold. Notably, DDIM outperforms the 250-step DDPM in the UCF101 task for both Latte and Skip-DiT .
Method UCF101 FFS Sky Taichi
Latte 165.04 23.75 28.88 5.36 49.46 11.51 166.84 11.57
Skip-DiT 173.70 22.95 20.62 4.32 49.22 12.05 163.03 13.55
Skip-Cachen=2 165.60 22.73 23.55 4.49 51.13 12.66 167.54 13.89
DDIM+Skip-DiT 134.22 24.60 37.28 6.48 86.39 13.67 343.97 21.01
DDIM+Latte 146.78 23.06 39.10 6.47 78.38 13.73 321.97 21.86
Skip-Cachen=3 169.37 22.47 26.76 4.75 54.17 13.11 179.43 14.53
DDIM+Skip-DiT 139.52 24.71 39.20 6.49 90.62 13.80 328.47 21.33
[…]
Table 3: Text-to-video generation performance. Latency and speedup are calculated using a single A100 GPU. The notation $n=i$ represents caching high-level features in $x_{t}$ for reuse during the inference of $x_{t-1},x_{t-2},\dots,x_{t-n+1}$. Skip-Cache 65% and Skip-Cache 75% correspond to performing caching at 65% and 75% of the timesteps, respectively. Baseline DiT models (without caching) are highlighted in grey, while the best metrics are emphasized in bold.
Method VBench(%) $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FLOPs (T) latency (s) speedup
Latte 76.14 1587.25 27.11 1.00$\times$
@@ 4.2 Main Results
PAB235 73.79   ($\downarrow$2.35) 19.18 0.27 0.66 1288.08 23.24 1.24$\times$
PAB347 72.08   ($\downarrow$4.06) 18.20 0.32 0.63 1239.35 22.23 1.29$\times$
PAB469 71.64   ($\downarrow$4.50) 17.40 0.35 0.60 1210.11 21.60 1.33$\times$
Skip-DiT 75.60 1648.13 28.72 1.00$\times$
Skip-Cache 75%
@@
Skip-Cache 65%
$n=2$ 75.51  ($\downarrow$0.09) 29.52 0.06 0.89 1127.83 19.28 1.49$\times$
$n=3$ 75.26  ($\downarrow$0.34) 27.46 0.09 0.85 974.80 16.67 1.72$\times$
$n=4$ 74.73  ($\downarrow$0.87) 25.97 0.13 0.81 882.98 15.12 1.90$\times$
@@
Table 4: Text-to-image generation performance. Latency and speedup are calculated on one A100 GPU. $n_{i}$ indicates caching the high-level feature in $x_{t}$ for reuse during the inference of $x_{t-1},x_{t-2},\dots,x_{t-n+1}$. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.

### v4

@@ 3.3 Integerating Long-Skip-Connection into DiT
To detail the DiT with Long-Skip-Connections(LSCs), we connect the shallow blocks to deep blocks with LSCs in DiT, named as Skip-DiT. Let $\bm{x}$ denote the input noise embedding, and $\bm{x}^{\prime}_{l}$ represents the output at the $l$-th layer of Skip-DiT. The architecture consists of $L$ sequential DiT blocks with LSCs. Each DiT block $\mathcal{F}_{\text{DiT}}^{l}$ at block $l$ processes the features as $\bm{x}^{\prime}=\mathcal{F}_{\text{DiT}}^{l}(\bm{x})$. The $i$-th skip branch ($i\in\{1\ldots L//2\}$) connects $i$-th block to ($L+1-i$)-th block, which can be denoted as $\mathcal{F}_{\text{Skip}}^{i}(\cdot,\cdot)$. Given output $\bm{x}^{\prime}_{i}$ from the start of the skip branch and $\bm{x}^{\prime}_{l}$ from the previous layer, the skip branch aggregates them to the input to $l$-th block as:
@@
where $\oplus$ denotes concatenation, Norm represents the layer normalization, and Linear is a fully-connected layer. Each skip connection creates a shortcut path that helps preserve and process information from earlier layers, enabling better gradient flow and feature reuse throughout the network.
3.4 Spectral Analysis of Skip-DiT
Consider an ideal denoising diffusion transformer $M$ with $L$ identical blocks, where the denoising capability induces the following fundamental properties:
  • Noise Reduction Invariance: Each transformer block $\mathcal{T}_{l}$ strictly reduces noise magnitude of input $h$. This directly implies the Jacobian spectral norm ($\sigma_{max}$) constraint:
    $\sigma_{\max}\left(\frac{\partial\mathcal{T}_{l}}{\partial h}\right)\triangleq\gamma_{l}<1$ (5)
  • Transformer Blocks Homogeneity: Identical noise reduction ratio across layers $\gamma_{l}=\gamma,\quad\forall l\in\{1,...,L\}$. Thus, the complete model satisfies
    $\sigma_{\max}(M)=\prod_{l=1}^{L}\gamma=\gamma^{L}\ll 1$ (6)
Theorem 1.
The spectral norm of the Jacobian matrix of DiT with Long-Skip-Connections is controlled more tightly than that of Vanilla DiT $M$, making the Skip-DiT model more robust, numerically stable, and capable of converging faster.
Proof.  Define layer transformations for $L/2<l\leq L$ and their Jacobian matrices, where $\mathcal{T}$ denotes the Transformer block and $10<\alpha<$ 1 denotes normlized concatetion:
$h_{l+1}^{\text{vanilla}}=\mathcal{T}(h_{l}^{\text{vanilla}}),h_{l+1}^{\text{skip}}=(1-\alpha)\cdot\mathcal{T}(h_{l}^{\text{skip}})+\alpha\cdot h_{L-l}^{\text{skip}}.$
The corresponding Jacobian matrices $J$ are:
$J_{l}^{\text{vanilla}}=\frac{\partial\mathcal{T}\left(h_{l}\right)}{\partial h_{l}},J_{l}^{\text{skip}}=(1-\alpha)\cdot\frac{\partial\mathcal{T}\left(h_{l}\right)}{\partial h_{l}}+\alpha\cdot\frac{\partial h_{L-l}}{\partial h_{l}}.$
Given $\gamma<1$ and $2l-L\geq 1$ for $l\geq L/2$, the layer wise spectral norms $\sigma_{max}$ of $M_{vanilla}$ and $M_{skip}$ staisfy:
$\sigma_{\max}(J_{l}^{\mathrm{skip}})$ $\leq(1-\alpha)\gamma+\alpha\gamma^{2l-L}<\gamma=\sigma_{\max}(J_{l}^{\mathrm{vanilla}})$
For the full model Jacobian $J=\prod_{l=1}^{L}J_{l}$, the $\sigma_{max}$ satisfy:
$\sigma_{\max}(J^{\mathrm{skip}})$ $<\prod_{l=0}^{L-1}\gamma=\gamma^{L}=\sigma_{\max}(J^{\mathrm{vanilla}}).$ (7)
This spectral norm reduction attenuates perturbation growth through $\|\Delta h_{L}\|\leq(\sigma_{\max}(J))^{L}\|\Delta h_{0}\|$, stabilizes gradient flow via $\|\nabla_{\theta}\mathcal{L}\|\leq\sigma_{\max}(J)\|\nabla_{h_{L}}\mathcal{L}\|$, and improves the Lipschitz constant $\mathrm{Lip}(M_{\mathrm{skip}})<\mathrm{Lip}(M_{\mathrm{vanilla}})$ with tight upper bound. This theoretical analysis aligns with our findings in Section 3.2, where Skip-DiT  are proven to be more feature stable and training efficient through comprehensive visualizations and preliminary experiments.
3.5 Cached Inference Mechanism
The feature stability and robustness in Skip-DiT enables effective reuse of intermediate features across timesteps during denoising. To quantify this advantage, we analyze the similarity of transformer block outputs with versus without caching in both vanilla DiT and Skip-DiT.
Caching Strategy for Skip-DiT
For Skip-DiT, we implement a hybrid caching strategy combining skip-connection-based block skipping from DeepCache [17], phased denoising strategy[13, 50], and errors accumulation guidance[11, 16]. As illustrated in Figure 3, our guidance mechanism is informed by a static error record $\mathcal{R}$, which is pre-computed by averaging historical prediction errors. To optimize the caching process, we also identify an optimal dynamic phase, $\mathcal{P}$, through static feature analysis. The choice $\mathcal{P}$ is substantiated by our ablation study in Table 5 and visualization in Figure 5. The complete inference procedure is formally outlined in Algorithm 1.
Visualizing Caching Stability
. As visualized in Figure 3, Skip-DiT  exhibits significantly higher feature similarity and smaller variance among examples compared to vanilla DiT, which employs FORA [30] and FasterDiffusion designed for DiT models. This demonstrates superior stability for cached inference. We attribute this improvement to the inherent feature stabilization and enhanced robustness. The latter can be theoretically grounded on $M$ defined in sec.3.4:
Theorem 2.
Under feature reuse with interval $\tau$, Skip-DiT achieves tighter perturbation bounds than Vanilla DiT, enabling larger allowable reuse intervals while maintaining error tolerance $\epsilon_{\mathrm{max}}$.
Proof.  Let $h_{t}$ be the reused feature with error $\epsilon_{t}$, $\delta$ the perturbation bound per step, $x$ the input, and $\theta$ the model parameters. Through the differential mean-value theorem and error propagation:
$\epsilon_{t+1}=\|h_{t+1}-h_{t}\|$ $\leq\mathrm{Lip}(M)\epsilon_{t}+\delta$
From Theorem 1, the Lipschitz constants and perturbation bounds satisfy: $\mathrm{Lip}(M_{\mathrm{skip}})<\mathrm{Lip}(M_{\mathrm{vanilla}})$ and $\delta_{\mathrm{skip}}<\delta_{\mathrm{vanilla}}$. For $T$-step feature reuse, the cumulative error is:
$\epsilon_{T}\leq\frac{\mathrm{Lip}(M)^{T}-1}{\mathrm{Lip}(M)-1}\delta$ (8)
Considering the conditions above, the maximal interval $\tau$ satisfies: $\epsilon_{\tau}^{skip}=\epsilon_{\tau}^{vanilla}<\epsilon_{max}$. Solving this yields $\tau_{\mathrm{skip}}>\tau_{\mathrm{vanilla}}$ under identical $\epsilon_{\mathrm{max}}$, which demonstrate better robustness and tolerance to error of Skip-DiT.
Algorithm 1 Skip-DiT  with Cached Inference
1:Initial state $\bm{x}^{t}$, model $\mathcal{F}$, cache interval $N$, global error $\epsilon$, static error records $\mathcal{R}(x_{t=1}^{T})$, dynamic Phase $\mathcal{P}$.
2:Generated sequence $\{\bm{x}^{t}\}_{t=1}^{T}$
3:Global Timestep $t$ (Full Inference):
4:$\bm{x^{\prime}}^{t}\leftarrow\mathcal{F}^{L-1}\circ\cdots\circ\mathcal{F}^{1}(\bm{x}^{t})$
5:Cache feature: $\mathcal{C}^{t}_{L-1}\leftarrow\bm{x^{\prime}}^{t}$
6:$\bm{x}^{t-1}=\mathcal{F}^{L}(\bm{x}^{t})$
7:for $k=1$ to $N-1$ do
8:     Local Timestep $t-k$ (Cached Inference):
9:     $\bm{x}^{\prime t-k}_{1}\leftarrow\mathcal{F}_{\text{DiT}}^{1}(\bm{x}^{t-k})$
10:     $\bm{x}^{t-k}\leftarrow\mathcal{F}_{\text{DiT}}^{L}\left(\mathcal{F}_{\text{Skip}}^{1}(\bm{x}^{\prime t-k}_{1},\mathcal{C}^{t}_{L-1})\right)$
11:     Maintain cache: $\mathcal{C}^{t}_{L-1}\leftarrow\mathcal{C}^{t}_{L-1}$
12:     Accumulate error: $\epsilon\leftarrow\epsilon+\mathcal{R}(t-k)$
13:     if $t-k-1\in\mathcal{P}\leavevmode\nobreak\ ||\leavevmode\nobreak\ \epsilon>\text{threshold}$ then
14:         $\epsilon\leftarrow 0$; Break
15:     end if
16:end for
17:$t\leftarrow t-k-1$;
18:Back to Full Inference until: $t<0$;
@@ 4 Experiments
Table 1: Class-to-video generation performance. The definition of the cache step $n$ follows that in Table 2. For the UCF101 task, the scheduler is 50-steps-DDIM, while for the remaining tasks is 250-steps-DDPM. Latency and speedup are calculated on one A100 GPU. The notation $n=i$ represents cache interval $N$ in Algorithm 1. We highlight Baseline DiT models in grey, and the best metrics in bold.
@@
Skip-DiT  with Cached Inference
original 141.30 23.78 20.62 4.32 49.21 11.92 163.03 13.55 290.05 10.02 1.00$\times$
@@ 4.1 Implementation Details
To start with, we add Long-Skip-Connections (LSCs) in DiT-XL [22] on the class-to-image task to prove the training efficiency and feature stability of  Skip-DiT. We employ Latte [18] as our base model in video generation and Hunyuan-DiT for text-to-image generation to further demonstrate the remarkable effectiveness of Skip-DiT  and its caching efficiency. Latte consists of spatial and temporal transformer blocks, suitable for both class-to-video and text-to-video tasks. Hunyuan-DiT [14], the first text-to-image DiT model with skip connections, is designed for text-to-image tasks. All DiT models are integrated with LSCs to evaluate the performance in both video and image generation tasks following the guidance in Section.3.3 and explore techniques for integrating LSCs into a pre-trained DiT model.
@@
For the class-to-image task, Skip-DiT is trained on 256$\times$256 ImageNet [7], following the dataset used by DiT-XL. In the class-to-video task, Skip-DiT  is trained on four datasets: FaceForensics [28], SkyTimelapse [42], UCF101 [34], and Taichi-HD [31]. Following the experimental settings in Latte, we extract 16-frame video clips from these datasets and resize all frames to a resolution of 256$\times$256. For the text-to-video task, the original Latte is trained on Webvid-10M [1] and Vimeo-2M [39], comprising approximately 330k text-video pairs in total. Considering that the resolution of Webvid-10M is lower than 512$\times$512, we only sample 330k text-video pairs from Vimeo-2M for training, with a resolution of 512$\times$512 and 16 frames.
@@
In the class-to-image task, the model is initialized randomly and aligns all the training settings with DiT-XL. The detailed training settings can be found in the Appendix. For class-to-video tasks, we train all instances of Skip-DiT  from scratch on all datasets. For text-to-video tasks, we propose a two-stage continual training strategy:
  • Skip-Connections training: Latte with LSC is initialized with the weights of the original Latte, and the weight of LSCs is initialized randomly. At this stage, only the LSCs are trained. It only takes one day on 8 H100 GPUs.
  • Overall training: With LSCs Fully trained, all other parameters are unfrozen to perform overall training. At this stage, Skip-DiT rapidly recovers its generation capability within two days and can generate content comparable to the original Latte with an extra three days of training.
All our training experiments are conducted on 8 H100 GPUs, video-image joint training strategy proposed by Ma et al. [18] is employed to enhance training stability.
4.2 Results on Video Generation Tasks
@@
Following previous works [53, 11, 16] and Latte, we evaluate text-to-video models using VBench [10], Peak Signal-to-Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [48], and Structural Similarity Index Measure (SSIM) [40]. VBench is a comprehensive benchmark suite comprising 16 evaluation dimensions. PSNR is a widely used metric for assessing the quality of image reconstruction, LPIPS measures feature distances extracted from pre-trained networks, and SSIM evaluates structural information differences. For class-to-image tasks, we evaluate the similarity between generated and real videos using Fréchet Video Distance (FVD) [38] and Fréchet Inception Distance (FID) [21], following the evaluation guidelines of StyleGAN-V [47].
Table 2: Text-to-video generation performance. Inference latency and flops are evaluated with 1 A100 GPU. The notation $n=i$ represents the cache interval $N$ in Algorithm 1. Baseline DiT models are highlighted in grey, and the best metrics are emphasized in bold.
Model Method VBench(%) $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FLOPs (T) Latency (s) Speedup
Latte original 76.14 1587.25 27.11 1.00$\times$
@@
PAB235 73.77   ($\downarrow$2.37) 19.18 0.27 0.66 1288.08 23.24 1.24$\times$
PAB347 71.11   ($\downarrow$5.03) 18.20 0.32 0.63 1239.35 22.23 1.29$\times$
PAB469 70.20   ($\downarrow$5.94) 17.40 0.35 0.60 1210.11 21.60 1.33$\times$
AdaCache 76.07   ($\downarrow$0.07) 22.23 0.19 0.78 934.85 16.95 1.60$\times$
TeaCache 76.03   ($\downarrow$0.11) 23.05 0.17 0.80 862.52 15.75 1.72$\times$
Skip-DiT original 75.60 1648.13 28.72 1.00$\times$
$65\%$ steps cached
$n=2$ 75.51  ($\downarrow$0.09) 29.52 0.06 0.89 1127.83 19.28 1.49$\times$
$n=3$ 75.26  ($\downarrow$0.34) 27.46 0.09 0.85 974.80 16.67 1.72$\times$
$n=4$ 74.73  ($\downarrow$0.87) 25.97 0.13 0.81 882.98 15.12 1.90$\times$
$75\%$ steps cached
@@
Class-to-video Generation
We compare the quantitative performance of Latte and Skip-DiT on four class-to-video tasks, as shown in Table 1. Skip-DiT consistently outperforms Latte in terms of FVD scores across all tasks while achieving comparable performance in FID scores, demonstrating its strong video generation capabilities. Furthermore, we observe that Skip-DiT  with cached inference significantly outperforms other caching methods across most metrics, incurring only an average loss of $2.37$ in the FVD score and $0.27$ in the FID score while achieving a 1.56$\times$ speedup.
@@
Table 2 presents evaluation results on text-to-video models. Videos are generated using the prompts of VBench [10], which is considered a more generalized benchmark [53, 10, 54]. Compared with the original Latte, Skip-DiT achieves a comparable VBench score (75.60 vs. 76.14) with only six days of continual pre-training on 330k training samples. To demonstrate the superiority of the cached inference of Skip-DiT , we evaluate two caching settings: caching at timesteps $700\rightarrow 50$ and $800\rightarrow 50$ (out of 1000 timesteps in total). In both settings, Skip-DiT achieves the highest speedup while maintaining superior scores in PSNR, LPIPS, and SSIM, with only a minor loss in VBench score.
4.3 Results on Image Generation Tasks
Table 3: Text-to-image generation performance. Latency and speedup are calculated on one A100 GPU. The notation $n=i$ represents cache interval $N$ in Algorithm 1. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.